# Behavior-Inspired Neural Networks for Relational Inference

### v2

@@ Abstract
From pedestrians to Kuramoto oscillators, interactions between agents govern how a multitude of dynamical systems evolve in space and time. Discovering how these agents relate to each other can improve our understanding of the often complex dynamics that underlie these systems. Recent works learn to categorize relationships between agents based on observations of their physical behavior. These approaches are limited in that the relationship categories are modelled as outcomes of categorical distribution, when in real world systems categories often intermingle and interact. In this work, we introduce a level of abstraction between the observable behavior of agents and the latent categories that determine their behavior. To do this, we learn a mapping from agent behavior to agent preferences for each latent category in a graph neural network. We integrate the physical proximity of agents and their preferences in a nonlinear opinion dynamics model which provides a mechanism to identify mutually exclusive latent categories, predict an agent’s evolution in time, and control an agent’s physical behavior. We demonstrate the utility of our model for learning interpretable categories, and its efficacy on long-horizon prediction across several benchmarks where we outperform existing methods.
@@ 1 Introduction
Multi-agent systems are found in domains as diverse as astronomy [1, 2], biology [3, 4, 5], physics [6, 7], and sports [8, 9]. Understanding how these systems evolve in time has the potential to provide insights useful for the discovery of unknown physics, the rules governing collective behavior, and engineering design. Predicting the evolution of complex systems is a fundamental challenge in the learning literature. Early black-box models determine future states from past states without regard for contextual information [10, 11]. More recent approaches improve upon these models by incorporating information about environmental conditions and the behavior of other agents [12, 13, 14, 15, 16]. While incorporating contextual information has led to more powerful trajectory prediction models, the opacity of these models limits our ability to leverage them to better understand the role of inter-agent relationships.
Recent work in relational inference attempts to address this limitation by explicitly modelling inter-agent relationships [17, 18, 19, 20, 21, 22]. Discovering inter-agent relationships is challenging, however, since ground truth labels are typically unavailable, and the relevant relationships may be unknown at design time. Graph neural network (GNN) based approaches such as  [19, 20, 21, 22] model inter-agent relationship categories with a categorical variable. In their seminal work, the authors of [19] learn a latent representation of inter-agent relationships in a trajectory prediction pipeline. The authors of [20] improve on this approach by modeling inter-agent relationships as time varying and the authors of [21] improve model expressivity by representing node features as a hypergraph. While these methods are able to predict future observations of a number of systems, the underlying assumption that inter-agent relationships are determined by a single relationship category diverges from what we observe in the real world.
In contrast to these models, opinion dynamics models assume agents have preferences for multiple categories, and that these preferences evolve in time. The nonlinear opinion dynamics (NOD) model introduced in [23], has been used to model a variety of societal systems [24, 25, 26, 27, 28, 29, 30, 31]. The nonlinear nature of this model introduces a bifurcation which allows preferences to quickly and flexibly respond to changes in environmental inputs. A limitation of this method, however, is that the relationship between agent preferences for a set of categories and their observable behavior must be known a priori.
@@ 2 Related Work
In contrast to traditional control approaches which tune a set of model parameters from all available trajectories (i.e., sequences of physical positions) [32, 33, 34, 35], early deep learning approaches learn to map a sequence of input states to future states directly [11]. Later efforts incorporated external influences (e.g., behavior of other agents, environmental conditions) [12, 13, 14, 15, 16] and temporal dependencies [16, 36, 37, 38, 39], and recent work has used graph based methods to model multi-agent dynamics [40, 41]. While these methods are able to predict future system states , they do not predict which agents interact or how they interact, a limitation which motivates relational inference.
Relational inference.
The goal of relational inference is to infer inter-agent relationships in a multi-agent system [17, 19, 20, 21, 22]. This task is challenging since, in general, the relationships between agents is unknown. Approaches to this challenge include: CommNet [17] which learns a continuous communication protocol for fully cooperative multi-agent tasks; Neural Relational Inference [19] which learns time invariant inter-agent relationships; Dynamic Neural Relational Inference [20] which learns time varying inter-agent relationships; GroupNet [21] which uses a multiscale hypergraph representation of the input for improved expressivity; and EquiMotion [22] which incorporates an equivariance constraint to ensure the dynamics transform predictably to Euclidean shifts of the input data. Our work differs from these approaches in our representation of relationship categories. Our categories exist in the space of agent preferences rather than the physical space, and our categories are flexible and interacting, rather than mutually exclusive and independent.
Consensus dynamics.
@@
In control and robotics, consensus dynamics [42] have been used in a myriad of settings to model the dynamics and control the behavior of multi-agent systems. In [43], the authors propose a linear model for prediction and control of multi-agent systems. In [44], the authors use a linear consensus dynamics model for coordinated surveying with underwater gliders. In [45], the authors use the consensus dynamics framework to develop rectilinear and circular consensus control laws for multi-vehicles systems. In [46], the authors use consensus dynamics to improve data collection in a mobile sensor network. In [47], the authors use consensus dynamics to understand the robustness of starling flocking behavior. Even with this breadth of application, there are several draw backs to a linear model of opinion formation; specifically, the naive implementation results in dynamics that only yields consensus [48, 49] and the formation of opinions in response to inputs is slow.
Nonlinear opinion dynamics.
@@
The noted short comings of linear consensus dynamics models are resolved in the nonlinear opinion dynamics model proposed in [23]. As the model is nonlinear, the dynamics result in a bifurcation and opinions can evolve to dissensus quickly. Nonlinear opinion dynamics have been used to model a variety of systems. In [24, 25, 26, 27] nonlinear opinion dynamics model information spread in settings such as political polarisation. In [28], nonlinear opinion dynamics is used to resolve deadlock, and in [29], it is used for collision avoidance. In [30, 31], nonlinear opinion dynamics is used for task switching in robotic swarms in trash collecting robot teams. In contrast to these works, where the model has direct access agent opinions (i.e., preferences), we learn a mapping between physical states and agent preferences.
@@ 3 Background
Nonlinear dynamics differ from linear dynamics in that they are able to exhibit bifurcations [50, 51, 52]. Nonlinear opinion dynamics [53] leverage bifurcations for fast and flexible decision making even with weak input signals [23]. Our BINN model integrates a nonlinear opinion dynamics inductive bias to interpretably and flexibly model agent behavior.
@@ 3.1 Nonlinear opinion dynamics
Consider a multi-agent system of $\mathcal{N}_{\mathrm{a}}\in\mathbb{N}$ agents, each possessing real-valued preferences about $\mathcal{N}_{\mathrm{o}}\in\mathbb{N}$ categories. Each category represents a belief or desire (e.g., take a physical action, undertake a task), and an agent’s preference for a category can be positive, neutral, or negative (the preference magnitude corresponds to preference strength). Changes to an agent’s preference for a category depend on extrinsic and intrinsic parameters. Concretely, the changes to agent $i$’s preference for category $j$ can be determined by the following nonlinear differential equation proposed in [23]
{dmath}˙z
_ij = -d_ijz_ij + S(u_i(α_ijz_ij + ∑_k=1k≠i^N_aa^a_ikz_kj + ∑_l=1l≠j^N_oa^o_jlz_il + ∑_k=1k≠i^N_a∑_l=1l≠j^N_o a^a_ika^o_jlz_kl)) + b_ij.
@@
The parameters $a^{a}_{jk}$, $a^{o}_{jl}$, and $b_{ij}$ are extrinsic to the agent. The communication matrix $[a^{a}_{ik}]$ describes the communication strength between agent $i$ and agent $k$, the belief matrix $[a^{o}_{jl}]$ describes the correlation of preferences for category $j$ and category $l$, and the environmental input $b_{ij}$ describes the impact of the environment on an agent’s the preference for category $j$. The saturating function $\mathcal{S}$ is selected so that $S\left(0\right)=0$, $S^{\prime}\left(0\right)=1$, and $S^{\prime\prime\prime}\left(0\right)\neq 0$ [23]. We use a hyperbolic tangent for the saturation function in our experiments.
For a given communication matrix $[a^{o}_{jl}]$ and belief matrix $[a^{o}_{jl}]$, the sensitivity of preference formation is determined by an agent’s intrinsic parameters $d_{ij}$, $u_{i}$, and $\alpha_{ij}$; and the equilibrium value of each preference is a function of the environmental input $b_{ij}$ (see Figure 2).
@@
Figure 2: Sensitivity to environmental inputs $b_{ij}$. Solid lines represent stable equilibria and dotted line represents unstable equilibria. (a) We show the pitchfork bifurcation characteristic to equation 3.1. The number, location, and stability of equilibria changes with the attention parameter $u_{i}$. (b) For attention $u_{i}<u^{*}$ preferences change linearly with environmental inputs. (c) For attention $u_{i}>u^{*}$ preferences change rapidly in response to environmental inputs with hysteresis encoding memory of previous states.
@@
Mutually exclusive categories.
In the nonlinear opinion dynamics framework, mutually exclusive categories are negatively correlated. Consider, for example, a system with two categories. The categories are mutually exclusive if $a^{o}_{12},a^{o}_{21}\leq 0$. Since $z_{i1}=-cz_{i2}$, we can decouple the dynamics for $z_{i1}$ and $z_{i2}$ and only model the preference for one category. Equation (3.1) therefore reduces to
@@
Additional details are provided in Appendix C.1. Mutually exclusive categories provides a systematic way to identify opportunities for dimensionality reduction.
@@ 4 Method
In this section we present our BINN model for relational inference (see Figure 3). Given trajectories of a multi-agent system, our goal is to predict agent behavior in an interpretable way by inferring the intrinsic and extrinsic characteristics of agents as described by the nonlinear opinion dynamics model (see Section 3).
We define a multi-agent system as a system of $\mathcal{N}_{\mathrm{a}}\in\mathbb{N}$ agents, each with real-valued preference for a set of $\mathcal{N}_{\mathrm{o}}\in\mathbb{N}$ categories. We learn a mapping between the observed behavior of agents and their preferences for a set of latent categories using a graph neural network; and the parameters of a nonlinear opinion dynamics model which determine the evolution of agent preferences.
We model the evolution of agent preference by the dynamical equation presented in Equation 3.1. With this formulation, we can control agent preferences by varying the environmental input as shown in Figure 5.
We train our model using $N$ trajectories of $T$ observations, where each observation has dimension $d$. For a given trajectory, we denote the observed state of agent $i$ at time $t$ by $\mathbf{x}_{i,t}\in\mathbb{R}^{d}$, and the preferences of agent $i$ at time $t$ by $\mathbf{z}_{i,t}\in\mathbb{R}^{\mathcal{N}_{o}}$.
@@ 4.1 Encoder
We use separate encoding networks to learn the mappings from agent states to agent preferences and from agent states to environmental inputs. We refer to the first of these encoders as the preference encoder $E_{z}$, and the latter as the environmental input encoder $E_{b}$. Each encoder takes state observations of the multi-agent system at time $t$ as input and processes them in an MPNN [54]. The multi-agent system is represented as a fully-connected graph with node values determined by the physical state of each agent.
Formally, the preference encoder $E_{z}$ performs the following message passing functions for agent $i$ at timestep $t$:

### v3

@@ Abstract
From pedestrians to Kuramoto oscillators, interactions between agents govern how dynamical systems evolve in space and time. Discovering how these agents relate to each other has the potential to improve our understanding of the often complex dynamics that underlie these systems. Recent works learn to categorize relationships between agents based on observations of their physical behavior. These approaches model relationship categories as outcomes of a categorical distribution which is limiting and contrary to real-world systems, where relationship categories often intermingle and interact. In this work, we introduce a level of abstraction between the observable behavior of agents and the latent categories that determine their behavior. To do this, we learn a mapping from agent observations to agent preferences for a set of latent categories. The learned preferences and inter-agent proximity are integrated in a nonlinear opinion dynamics model, which allows us to naturally identify mutually exclusive categories, predict an agent’s evolution in time, and control an agent’s behavior. Through extensive experiments, we demonstrate the utility of our model for learning interpretable categories, and the efficacy of our model for long-horizon trajectory prediction.
@@ 1 Introduction
Multi-agent systems are found in domains as diverse as astronomy (Villanueva-Domingo et al., 2022; Lemos et al., 2023), biology (Fiorelli et al., 2006; Seeley et al., 2012; Young et al., 2013), physics (Gull et al., 2013; Browaeys & Lahaye, 2020), and sports (Hauri et al., 2021; Wang et al., 2024). Understanding how these systems evolve in time has the potential to provide insights useful for the discovery of unknown physics, the rules governing collective behavior, and engineering design. Predicting the evolution of complex systems is a fundamental challenge in the learning literature. Early black-box models determine future states from past states without regard for contextual information (Hochreiter & Schmidhuber, 1997; Sutskever et al., 2014). More recent approaches improve upon these models by incorporating information about environmental conditions and the behavior of other agents (Alahi et al., 2016; Gupta et al., 2018; Casas et al., 2018; Deo & Trivedi, 2018; Sadeghian et al., 2019). While incorporating contextual information has led to more powerful trajectory prediction models, the opacity of these models limits our ability to leverage them to better understand the role of inter-agent relationships.
Recent work in relational inference attempts to address this limitation by explicitly modeling inter-agent relationships (Sukhbaatar et al., 2016; Santoro et al., 2017; Kipf et al., 2018; Graber & Schwing, 2020; Xu et al., 2022, 2023). Discovering inter-agent relationships is challenging, however, since ground truth labels are typically unavailable, and the relevant relationships may be unknown at design time. Graph neural network (GNN) based approaches such as  Kipf et al. (2018); Graber & Schwing (2020); Xu et al. (2022, 2023) model inter-agent relationship categories with a categorical variable. In their seminal work, the authors of Kipf et al. (2018) learn a latent representation of inter-agent relationships in a trajectory prediction pipeline. The authors of Graber & Schwing (2020) improve on this approach by modeling inter-agent relationships as time varying and the authors of Xu et al. (2022) improve model expressivity by representing node features as a hypergraph. While these methods are able to predict future observations of a number of systems, the underlying assumption that inter-agent relationships are determined by a single relationship category diverges from what we observe in the real world.
In contrast to these models, opinion dynamics models assume agents have preferences for multiple categories, and that these preferences evolve in time. The nonlinear opinion dynamics (NOD) model introduced in Leonard et al. (2024), has been used to model a variety of societal systems (Das et al., 2014; Rossi & Frasca, 2020; Leonard et al., 2021; Franci et al., 2021; Cathcart et al., 2022; Bizyaeva et al., 2022a, 2024; Hu et al., 2023; Ordorica Arango et al., 2024; Wang et al., 2025). The nonlinear nature of this model introduces a bifurcation which allows preferences to quickly and flexibly respond to changes in environmental inputs. A limitation of this method, however, is that the relationship between agent preferences for a set of categories and their observable behavior must be known a priori.
@@ 2 Related Work
Trajectory prediction. In contrast to traditional control approaches which tune a set of model parameters from all available trajectories (Brunton et al., 2016; Galioto & Gorodetsky, 2020; Paredes & Bernstein, 2021, 2024; Paredes et al., 2024; Richards et al., 2024), early deep learning approaches learn to map a sequence of input states to future states directly (Sutskever et al., 2014). Later efforts incorporated external influences (e.g., behavior of other agents, environmental conditions) (Alahi et al., 2016; Gupta et al., 2018; Casas et al., 2018; Deo & Trivedi, 2018; Sadeghian et al., 2019), physical priors (Greydanus et al., 2019; Mason et al., 2022, 2023; Allen-Blanchette, 2024), and temporal dependencies (Vemula et al., 2018; Sadeghian et al., 2019; Mangalam et al., 2020; Yuan et al., 2021; Giuliari et al., 2021; Rezaei & Dieng, 2024), and recent work has used graph based methods to model multi-agent dynamics (Yu et al., 2020; Gao et al., 2020). While these methods can predict future system states, they do not predict which agents interact or how they interact, a limitation which motivates relational inference.
Relational inference. The goal of relational inference is to infer inter-agent relationships in a multi-agent system. This task is challenging since, in general, the relationships between individual agents are unobservable. Early works (Sukhbaatar et al., 2016; Foerster et al., 2016) focus on learning the communication protocols for multi-agent systems. Neural Relational Inference (Kipf et al., 2018) proposed a variational autoencoding (Kingma et al., 2013) graph neural network framework for learning time-invariant relationships between individual agents. Further developments focused on increasing the expressivity and applicability of this framework by incorporating factorized graphs to model different types of interactions (Webb et al., 2019); dynamic encoders to model time-varying inter-agent relations Graber & Schwing (2020); edge-to-edge message passing for more efficient information sharing (Chen et al., 2021); shared decoder conditioned on edge types to learn a mapping for future states (Löwe et al., 2022); hypergraphs to accommodate interactions of different spatial scales (Xu et al., 2022); and Euclidean transformation equivariance to improve generalization to varying scenes Xu et al. (2023). Other works leverage different architectures to infer inter-agent relations such as meta-learning to map inputs to edge and node values (Alet et al., 2019); reservoir computing to increase the efficiency of relational inference (Wang et al., 2023); partial correlates of latent representations to infer connections (Wang & Pang, 2024); and diffusion models to reconstruct missing connections (Zheng et al., 2024). Our work differs from these approaches in our representation of relationship categories. Our categories exist in a space of agent preferences rather than the physical space, and our categories are flexible and interacting, rather than mutually exclusive and independent.
@@
Consensus dynamics. In control and robotics, consensus dynamics (Bullo, 2018) have been used in a myriad of settings to model the dynamics and control the behavior of multi-agent systems. In Levine et al. (2000), the authors propose a linear model for prediction and control of multi-agent systems. In Leonard et al. (2010), the authors use a linear consensus dynamics model for coordinated surveying with underwater gliders. In Justh & Krishnaprasad (2005), the authors use the consensus dynamics framework to develop rectilinear and circular consensus control laws for multi-vehicles systems. In Leonard et al. (2007), the authors use consensus dynamics to improve data collection in mobile sensor networks. In Ballerini et al. (2008), the authors use consensus dynamics to understand the robustness of starling flocking behavior. Even with this breadth of application, there are drawbacks to a linear model of opinion formation; specifically, the naive implementation results in dynamics that only yields consensus (Altafini, 2012; Dandekar et al., 2013) and the formation of opinions in response to inputs is slow.
@@
Nonlinear opinion dynamics. The noted short comings of linear consensus dynamics models are resolved in the nonlinear opinion dynamics model proposed in Leonard et al. (2024). As the model is nonlinear, the dynamics result in a bifurcation and opinions can evolve to dissensus quickly. Nonlinear opinion dynamics have been used to model a variety of systems. In Das et al. (2014); Rossi & Frasca (2020); Leonard et al. (2021); Franci et al. (2021) nonlinear opinion dynamics model information spread in settings such as political polarisation. In Hu et al. (2023), nonlinear opinion dynamics is used to resolve deadlock, and in Cathcart et al. (2022), it is used for collision avoidance in human-robotic systems. In Bizyaeva et al. (2024); Ordorica Arango et al. (2024), nonlinear opinion dynamics is used for modeling risk in epidemic models. In contrast to these works, where the model has direct access agent opinions (i.e., preferences), we learn a mapping between physical states and agent preferences.
@@ 3 Background
Nonlinear dynamics differ from linear dynamics in that they are able to exhibit bifurcations (Golubitsky et al., 2012; Guckenheimer & Holmes, 2013; Strogatz, 2018). Nonlinear opinion dynamics (Bizyaeva et al., 2022b) leverage bifurcations for fast and flexible decision making even with weak input signals (Leonard et al., 2024). Our BINN model integrates a nonlinear opinion dynamics inductive bias to interpretably and flexibly model agent behavior.
@@ 3.1 Nonlinear opinion dynamics
Consider a multi-agent system of $\mathcal{N}_{\mathrm{a}}\in\mathbb{N}$ agents, each possessing real-valued preferences about $\mathcal{N}_{\mathrm{o}}\in\mathbb{N}$ categories. Each category represents a belief or desire (e.g., take a physical action, undertake a task), and an agent’s preference for a category can be positive, neutral, or negative (the preference magnitude corresponds to preference strength). Changes to an agent’s preference for a category depend on extrinsic and intrinsic parameters. Concretely, the changes to agent $i$’s preference for category $j$ can be determined by the differential equation proposed in Leonard et al. (2024)
$\dot{z}_{ij}=-d_{ij}z_{ij}+\mathcal{S}\Bigg{(}u_{i}\Bigg{(}\alpha_{ij}z_{ij}+\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}a^{\mathrm{a}}_{ik}z_{kj}\\ +\sum_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{\mathcal{N}_{\mathrm{o}}}a^{\mathrm{o}}_{jl}z_{il}+\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}\sum_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{\mathcal{N}_{\mathrm{o}}}a^{\mathrm{a}}_{ik}a^{\mathrm{o}}_{jl}z_{kl}\Bigg{)}\Bigg{)}+b_{ij}.$ (1)
@@
The parameters $a^{a}_{jk}$, $a^{o}_{jl}$, and $b_{ij}$ are extrinsic to the agent. The communication matrix $[a^{a}_{ik}]$ describes the communication strength between agent $i$ and agent $k$, the belief matrix $[a^{o}_{jl}]$ describes the correlation of preferences for category $j$ and category $l$, and the environmental input $b_{ij}$ describes the impact of the environment on agent $i$’s preference for category $j$. The saturating function $\mathcal{S}$ is selected such that $S\left(0\right)=0$, $S^{\prime}\left(0\right)=1$, and $S^{\prime\prime\prime}\left(0\right)\neq 0$ (Leonard et al., 2024). We use $\operatorname{tanh}\left(\cdot\right)$ for the saturation function in our experiments.
For a given communication matrix $[a^{o}_{jl}]$ and belief matrix $[a^{o}_{jl}]$, the sensitivity of preference formation is determined by an agent’s intrinsic parameters $d_{ij}$, $u_{i}$, and $\alpha_{ij}$, and the equilibrium value of each preference is a function of the environmental input $b_{ij}$ (see Figure 2).
@@
Figure 2: Sensitivity to environmental inputs $b_{ij}$. Solid lines represent stable equilibria and dotted line represents unstable equilibria. (a) We show the pitchfork bifurcation characteristic to Equation (1). The number, location, and stability of equilibria changes with the attention parameter $u_{i}$. (b) For attention $u_{i}<u^{*}$ preferences change linearly with environmental inputs. (c) For attention $u_{i}>u^{*}$ preferences change rapidly in response to environmental inputs with hysteresis encoding memory of previous states.
@@
Mutually exclusive categories. In the nonlinear opinion dynamics framework, the presence of mutually exclusive categories simplifies the model. For example, consider the case of $\mathcal{N}_{a}$ agents and $\mathcal{N}_{o}=2$ categories. The categories are said to be mutually exclusive if $a^{o}_{12},a^{o}_{21}\leq 0$ and $z_{i1}=-cz_{i2}$ (i.e. a positive preference for one category implies a negative preference for the other). In this setting, the dynamics for $z_{i1}$ and $z_{i2}$ decouple, and Equation (1) reduces to
@@
In our model, we leverage the presence of mutually exclusive categories to systematically identify opportunities for dimensionality reduction. Additional details are provided in Appendix 11.
@@ 4 Method
In this section we present our BINN model for relational inference (see Figure 3). Given trajectories of a multi-agent system, our goal is to predict agent behavior in an interpretable way by inferring the intrinsic and extrinsic characteristics of agents as described by the nonlinear opinion dynamics model (see Section 3 and Appendix D).
We define a multi-agent system as a system of $\mathcal{N}_{\mathrm{a}}\in\mathbb{N}$ agents, each with real-valued preferences for a set of $\mathcal{N}_{\mathrm{o}}\in\mathbb{N}$ categories. We learn a mapping between the observed behavior of agents and their preferences for a set of latent categories using a graph neural network, and the parameters of a nonlinear opinion dynamics model which determine the evolution of agent preferences.
We model the evolution of agent preference by the dynamical equation presented in Equation (1). With this formulation, we can control agent preferences by varying the environmental input as shown in Figure 5.
We train our model using $N$ trajectories of $T$ observations, where each observation has dimension $d$. For a given trajectory, we denote the observed state of agent $i$ at time $t$ by the concatenation of position and velocity $\mathbf{x}_{i,t}=\left[\mathbf{p}_{i,t},\mathbf{v}_{i,t}\right]\in\mathbb{R}^{d}$, and the preferences of agent $i$ at time $t$ by $\mathbf{z}_{i,t}\in\mathbb{R}^{\mathcal{N}_{o}}$.
@@ 4.1 Encoder
We use separate encoding networks to learn the mappings from agent states to agent preferences and from agent states to environmental inputs. We refer to the first of these encoders as the preference encoder $E_{z}$, and the latter as the environmental input encoder $E_{b}$. Each encoder takes state observations of the multi-agent system at time $t$ as input and processes them in an message passing neural network (MPNN) (Gilmer et al., 2017). The multi-agent system is represented as a fully-connected graph with node values determined by the physical state of each agent.
Our preference encoder $E_{z}$ performs the following message passing functions for agent $i$ at timestep $t$: